# ACTIVE: Continuous Similarity Search for Vessel Trajectories

### v2

@@ 3.2. Segment Partitioning
(1) $F[j]=\textit{min}\{F[j-l]+\mathcal{M}_{\textit{area}}(p^{j-l+1},p^{j})|l_{\textit{min}}\leq l\leq l_{\textit{max}}\},$
@@
where $\mathcal{M}_{\textit{area}}(p^{j-l+1},p^{j})$ is the MBR area of segment $\langle p^{j-l+1},...,p^{j}\rangle$. Parameter $l_{\textit{min}}$ plays a critical role in controlling the number of segments, which directly impacts the construction of the R-tree-based index structure. By reducing the number of segments, the efficiency of tree indexing and query processing is improved. Meanwhile, $l_{\textit{max}}$ defines an upper bound on the time complexity of pruning operations. This ensures that the computational cost remains manageable, even as the trajectory data becomes more complex. By balancing segment length constraints and MBR optimization, the partitioning process enhances spatial query performance and overall system efficiency in managing large-scale trajectory data.
4. Similarity Measure
4.1. Case Analysis
Figure 4. Examples of gaps between existing similarity measures and real needs
Case 1: Trend Neglect. Traditional trajectory similarity measures focus on minimizing geometric or temporal differences without considering the directional movement. In Figure 4(a), the distance between trajectories $T_{1}$ and $T_{o}$ is computed to be the same as that between $T_{2}$ and $T_{o}$ under existing measures. However, in many real-world applications, such as predictive modeling or motion forecasting, the trajectory trend plays a critical role in determining similarity. For example, although $T_{2}$ is farther from $T_{o}$ at earlier timestamps, its trend more closely aligns with $T_{o}$’s current movement direction, suggesting that $T_{2}$ is more likely to converge with $T_{o}$ in the near future. This indicates that ignoring the trend can lead to suboptimal or misleading results, particularly in scenarios where the future behavior of trajectories is of primary interest.
Case 2: Segment Relevance. Conventional trajectory similarity measures assume that all segments or points along a trajectory contribute equally to the overall similarity computation. This assumption overlooks the fact that, in many downstream tasks, certain segments of a trajectory may carry more relevance than others. For example, in Figure 4(b), given a trajectory $T_{o}$ and a set of historical trajectories $T_{1},T_{2}$}, only the segment highlighted in the red box is of primary interest for comparison. Ignoring the varying importance of trajectory segments can lead to suboptimal decision-making in critical scenarios.
Case 3: Target Point Neglect. In many practical applications, such as logistics and real-time navigation, the target destination of a moving object is of paramount importance. Current trajectory similarity measures, however, are predominantly backward-looking, focusing on historical data while disregarding the significance of the target destination and the trajectory’s future course. For instance, as shown in Figure 4(c), while the historical paths of trajectories $T_{1}$ and $T_{2}$ are similar, $T_{1}$’ trend after the pivotal point of are more aligned with the projected path of $T_{o}$ when the target destination is taken into consideration. Incorporating the target point into trajectory similarity measures would not only enhance the accuracy of real-time trajectory search but also enable more reliable predictions of future movements in applications such as autonomous navigation, shipping logistics, and traffic management.
4.2. Object-Trajectory Real-time Distance
Definition 0 (Object-Trajectory Real-time Distance).
Given a moving object $o$ at timestamp $t$ with information $(p_{o}^{c},T_{o})$ and a trajectory $T_{i}=\langle p_{i}^{1},p_{i}^{2},\dots,p_{i}^{n}\rangle$, the Object-Trajectory Real-time Distance between $o$ and $T_{i}$ is defined as follows:
(2) $\mathcal{D}(o,T_{i})=\alpha\cdot\mathcal{D}_{H}(T_{o},T^{\prime}_{i})+(1-\alpha)\cdot\mathcal{D}_{T}(p_{o}^{d},T_{i}^{L})$
[…]
(4) $\mathcal{D}_{T}(p_{o}^{d},T_{i}^{L})=d(p_{o}^{d},L(p_{i}^{x},p_{i}^{n}))$
Here, $d(p_{o}^{*},p_{i}^{*})$ is the euclidean distance between $p_{o}^{*}$ and $p_{i}^{*}$, where $p_{o}^{*}$ and $p_{i}^{*}$ are last points in $T_{o}$ and $T^{\prime}_{i}$. $\theta\in[0,1]$ is the decay factor. $x=\text{argmin}_{0<j\leq n}d(p_{o}^{c},p_{i}^{j})$ is the ID of the pivotal point in $T_{i}$. $T_{o}^{*}$ and $T^{\prime}_{i}$$^{*}$ are the segments of $T_{o}$ and $T^{\prime}_{i}$ with the last point removed. The distance can be flexibly customized by a trade-off parameter $\alpha\in[0,1]$ according to specific application needs. The process of OTRD computation is detailed in Section 5.2.
Figure 5. Example of OTRD measure
Example 0.
Figure 5 shows an example of OTRD. Given a trajectory $T_{1}=\langle(1,2),(3,3),(5,3),(6,5),(7,6),(9,7)\rangle$, a moving object $o$ at time $t$ with the target destination $p_{o}^{d}=(9,5)$ and $(p_{o}^{c},T_{o})$, where $p_{o}^{c}=(6,3)$ and $T_{o}=\langle(2,1),(4,2),(5,2),(6,3)\rangle$, the trade-off parameter $\alpha=0.9$, and the decay factor $\theta=0.9$, we have the pivotal point of $T_{1}$, $p_{1}^{x}=p_{1}^{3}=(5,3)$, and two parts of $T_{1}$, $T^{\prime}_{1}=\langle(1,2),(3,3),(5,3)\rangle$ and $T_{1}^{L}=L((5,3),(9,7))$. The HTD and TTD can be computed accordingly, i.e., $\mathcal{D}_{H}(T_{o},T^{\prime}_{1})=1.14$ and $\mathcal{D}_{T}(p_{o}^{d},T_{1}^{L})=1.41$. Therefore, the OTRD between $o$ and $T_{1}$ is $\mathcal{D}(o,T_{1})=0.9*1.14+0.1*1.41=1.167$.
5. Search Algorithm
5.1. Speed-up Strategies
@@
(5) $\mathcal{D}_{s}(p,\mathcal{M}(s))=\text{min}\{d(p,v_{l}^{u}),d(p,v_{l}^{b}),d(p,v_{r}^{u}),d(p,v_{r}^{b})\}.$
@@
When computing the HTD between the moving object $o$ at timestamp $t$ with information $o_{t}=(p_{o}^{c},T_{o})$ and a trajectory $T_{i}$, a segment $s_{i}$ can be pruned if
(6) $\theta^{c-j}\mathcal{D}_{s}(p_{o}^{j},\mathcal{M}(s_{i}))>\textit{dist}_{\textit{min}},$
@@
Given the current $k$-th smallest Object-Trajectory Real-time Distance $\textit{dist}_{k}$ between a moving object $o$ at timestamp $t$ with information $(p_{o}^{c},T_{o})$ and the computed historical trajectories, a trajectory $T_{i}$ can be pruned if
(7) $\mathcal{D}^{l}(p_{o}^{j},T^{\prime}_{i})\geq\frac{\textit{dist}_{k}-(1-\alpha)\mathcal{D}_{T}(p_{o}^{d},T_{i}^{L})}{\alpha\cdot\theta^{c-j}},$
@@
(8) $\mathcal{D}^{l}(p_{o}^{j},T^{\prime}_{i})=\text{min}\{d(p_{o}^{j},p_{i})|p_{i}\in T^{\prime}_{i}\}.$
Lemma 5.2.
Given a moving object $o$ at timestamp $t$ with information $o_{t}=(p_{o}^{c},T_{o})$ and a trajectory $T_{i}$, the Historical Trajectory Distance $\mathcal{D}_{H}(T_{o},T^{\prime}_{i})$ is no less than $d(p_{o}^{c},p_{i}^{x})$.
@@
According to Lemma 5.2, an early pruning strategy can be used before fully computing the distance between $o$ and $T_{i}$. $T_{i}$ can be pruned if $d(p_{o}^{c},p_{i}^{x})\geq(\textit{dist}_{k}-(1-\alpha)\cdot\mathcal{D}_{T}(p_{o}^{d},T_{i}^{L}))/\alpha$.
@@
Case 1: If the pivotal point ID in $T_{i}$ at $t$ is unchanged, then the TTD $\textit{dist}_{T}$ is unchanged and the HTD $\textit{dist}_{H}$ can be updated as follows.
(9) $\textit{dist}_{H}=\text{max}\{\theta\cdot\textit{dist}^{\prime}_{H},\text{min}\{d(p_{o}^{c},p_{i})|p_{i}\in T^{\prime}_{i}\}\}$
@@
Case 3: If the pivotal point ID in $T_{i}$ at $t$ is less than $x^{\prime}$, $\textit{dist}_{H}$, and $\textit{dist}_{T}$ can be computed according to Equations 3 and 4.
@@
However, this strategy may introduce errors to $\mathcal{D}_{H}(T_{o},T^{\prime}_{i})$ since some points in $T_{i}$ are ignored. We deduce the lower and upper error bounds as follows.
(10) $\mathcal{E}^{l}=\mathcal{D}_{H}(T_{o},T^{\prime}_{i})-d(p_{o}^{c},p_{i}^{x})$
(11) $\mathcal{E}^{u}=\text{max}\{\theta^{c-j}\cdot d(p_{o},p_{i}^{j})|p_{o}\in T_{o},p_{i}^{j}\in T^{\prime}_{i}(g)\}-\mathcal{D}_{H}(T_{o},T^{\prime}_{i})$
The lower error bound $\mathcal{E}^{l}$ can be easily proved according to Lemma 5.2. The upper error bound can be proved as follows.
@@
(13) $\gamma=|x-x^{\prime}|\mbox{ mod }g$
@@
5.2. OTRD Computation
The similarity computation between a moving object $o$ and a trajectory $T_{i}$ is performed in several key steps. First, the algorithm identifies the pivotal point $p_{i}^{x}$ of $T_{i}$ to the current position $p_{o}^{c}$ of $o$. Next, it computes the distance between $o$’ target destination $p_{o}^{d}$ and the line $L(p_{i}^{x},p_{i}^{n})$, where $p_{i}^{n}$ is the end point of $T_{i}$. Afterward, the distance between the moving object’s current trajectory $T_{o}$ and the sub-trajectory $T^{\prime}_{i}=\langle p_{i}^{1},...,p_{i}^{x}\rangle$ from $T_{i}$. To enhance the efficiency during this process, the speed-up strategies are employed to avoid unnecessary distance computations. Finally, the overall similarity is determined by summing these distances with a trade-off parameter $\alpha$, providing a comprehensive measure of how closely the moving object’s trajectory aligns with the historical one.
@@ 5.3. CSTS Algorithm
6. Experiments
6.1. Overall Experimental Settings
All experiments were conducted on a Linux server running Ubuntu with an Intel Xeon Icelake processor featuring 60 cores at 5.7 GHz and 240 GB of memory. The system is equipped with dual NVIDIA A10 GPUs, each providing 23 GB VRAM, for a total of 46 GB of GPU memory. All neural network models are implemented using PyTorch and trained on the GPU, and all index construction and search algorithms are implemented in Python (Code for CSTS [^1] https://github.com/hyLiu1994/CSTS/blob/main/README.md (Last access: 2025.1)).
Datasets. We evaluate the proposed methods on the following datasets. Dataset statistics are reported in Table 2.
  • RQ1. Evaluation on SVTI: How efficient is the construction process of SVTI?
  • RQ2. Evaluation on CSTS: How does the search algorithm CSTS perform in terms of efficiency and effectiveness?
  • RQ3. Ablation Study: What is the impact of the speed-up strategies on overall performance?
Table 2. Dataset statistics
@@ 6.1. Overall Experimental Settings
Table 3. Metrics and baselines
@@
RQ1 Construction Time (s), Index Size (MB) O-PQT, O-PRT, H/F-TrajI, H/F-DFT
RQ2 Query Time (ms), Hit (%) O-PQT, O-PRT, H/F-TrajI, H/F-DFT, TrajCL-TI, CLEAR-TI
@@
Metrics. The metrics for each task are listed in Table 3. To evaluate the efficiency of SVTI, we measure its construction time and index size. To evaluate the performance of CSTS, we select ten query instances, execute the selected queries ten times, and report the average. For effectiveness evaluation, we employ the hit rate ($\mathit{Hit}$):
(14) $\mathit{Hit}=\frac{\lvert\mathcal{T}_{f}\cap\mathcal{T}\rvert}{k},$
Baselines. We compare the proposed methods with several baseline methods, as outlined in Table 3. Specifically, we study SVTI and four traditional techniques with respect to index construction time and size. Furthermore, CSTS is compared with six traditional methods and two learning-based methods to evaluate its efficiency and effectiveness. The baseline methods are listed as follows:
@@
  • H-TrajI/F-TrajI: Hausdorff/Fréchet + Traj_Index (Xie et al., 2017). An R-Tree-based similar trajectory method using Hausdorff or Fréchet distance, where the index is built according to the centroid of each data trajectory’s minimum bounding rectangle.
  • H-DFT/F-DFT: Hausdorff/Fréchet + DFT (Xie et al., 2017). A Distributed Framework for Trajectory similarity search (DFT) using Hausdorff or Fréchet distance.
  • TrajCL-TI: TrajCL (Chang et al., 2023) + SVTI. A trajectory similarity search method based on SVTI, employing a graph-based long-term dependency model for similarity computation.
  • CLEAR-TI: CLEAR (Li et al., 2024a) + SVTI. A trajectory similarity search method using SVTI with a similarity computation technique based on contrastive representation learning.
@@
Parameters. The parameter settings are summarized in Table 4, with default values in bold. A description of each parameter is provided below:
  • $l_{\textit{min}}$ and $l_{\textit{max}}$: The minimum and maximum segment lengths.
@@
Table 4. Parameter settings
  Parameters Questions Settings
$l_{\textit{min}}$ RQ1 20, 25, 30, 35, 40
$l_{\textit{max}}$ RQ1 40, 45, 50, 55, 60
$\Lambda$ RQ1, RQ2 D-1, D-2, D-3, D-4, D-5
$l_{q}$ RQ2 10, 20, 30, 40, 50
$r$ RQ2 5, 10, 15, 20, 25
$k$ RQ2 30, 40, 50, 60, 70
$\theta$ RQ2 0.45, 0.5, 0.55, 0.6, 0.65
$\alpha$ RQ2 0.45, 0.5, 0.55, 0.6, 0.65
$g$ RQ3 1, 5, 10, 15, 20
6.2. Evaluation on CSTS

### v3

@@ IV-A Segment-based Trajectory Index
$F[j]=\textit{min}\{F[j-l]+\mathcal{M}_{\textit{area}}(p^{j-l+1},p^{j})|l_{\textit{min}}\leq l\leq l_{\textit{max}}\},$ (4)
@@
where $\mathcal{M}_{\textit{area}}(p^{j-l+1},p^{j})$ is the MBR area of segment $\langle p^{j-l+1},...,p^{j}\rangle$. Parameter $l_{\textit{min}}$ plays a critical role in controlling the number of segments, which directly impacts the construction of the R-tree-based index structure. By reducing the number of segments, the efficiency of tree indexing and query processing is improved. Meanwhile, $l_{\textit{max}}$ defines an upper bound on the time complexity of pruning operations. This ensures that the computational cost remains manageable, even as the trajectory data becomes more complex. By balancing segment length constraints and MBR optimization, the partitioning process enhances query performance and overall system efficiency in managing large-scale trajectory data.
2:Output: A segment set $\mathcal{S}$.
3:Initialize a segment set $\mathcal{S}$ $\leftarrow$ $\emptyset$
4:for $T_{i}=\langle p_{i}^{1},p_{i}^{2},...,p_{i}^{n}\rangle$ in $\mathcal{T}$ do
5:  Initialize a list for $T_{i}$’s segment set $S_{i}$ $\leftarrow$ $\emptyset$
6:  Initialize a list for the total MBR area $F$ $\leftarrow$ $\emptyset$
7:  for $j$ in $[1,n]$ do
8:   if $j\leq l_{\textit{min}}$ then $F[j]$ $\leftarrow$ 0
9:   else $F[j]$ $\leftarrow$ $\infty$
10:   for $l$ in $l_{\textit{min}},l_{\textit{max}}]$ do
11:     if $j-l\geq 0$ then
12:      $m_{l}$ $\leftarrow$ $\mathcal{M}_{area}$($p_{i}^{j-l+1}$, $p_{i}^{j}$)
13:      if jlj-l is 0 then $m$ $\leftarrow$ $m_{l}$
14:      else $m$ $\leftarrow$ $F[j-l]+m_{l}$
15:      if $F[j]>m$ then
16:        $F[j]$ $\leftarrow$ $m$
18:  $\mathcal{S}\leftarrow\mathcal{S}\cup S_{i}[n]$
19:return $\mathcal{S}$
Algorithm 1 outlines the segment partitioning process in detail. Initially, a segment set $\mathcal{S}$ is initialized (line 1). For each trajectory $T_{i}$, the algorithm identifies the optimal segmentation (lines 2–16). Specifically, a list is created for $T_{i}$’s segment set, where the $j$th item represents the current optimal segment set up to point $p^{j}$ (line 3). Additionally, another list is initialized to store the total MBR area, where the $j$th item represents the minimum total MBR area up to point $p^{j}$ (line 4). For each point $p^{j}$ in $T_{i}$, the minimum total MBR area up to $p^{j}$ is computed (lines 5–15). If $j$ is less than $l_{\textit{min}}$, the area is set to 0; otherwise, it is initialized to $\infty$ (lines 6–7). For each possible starting point $p^{j-l+1}$, where $l_{\textit{min}}\leq l\leq l_{\textit{max}}$, the MBR area of the segment $\langle p^{j-l+1},...,p^{j}\rangle$ is computed (lines 8–10). If the total MBR area is smaller than the current value, both the area and the segment set are updated (lines 11–15). Once all points in $T_{i}$ have been processed, the final segment set for $T_{i}$ is added to the overall segment set $\mathcal{S}$ (line 16). Ultimately, the algorithm returns the overall segment set $\mathcal{S}$ (line 17). The complexity of each trajectory $T_{i}$ partitioning is $O(n(l_{max}-l_{min}+1))$, where $n$ is the number of points in $T_{i}$.
IV-B Speed-up Strategies
@@
$\mathcal{D}_{s}(p,\mathcal{M}(s))=\text{min}\{d^{*}(p,v_{l}^{u}),d^{*}(p,v_{l}^{b}),d^{*}(p,v_{r}^{u}),d^{*}(p,v_{r}^{b})\},$ (5)
where $d^{*}(p,v)=\text{min}\{|p.x-v.x|,|p.y-v.y|\}$.
@@
When computing the HTD between the moving object $o$ at timestamp $t$ with $o_{t}=(p_{o}^{c},T_{o})$ and a trajectory $T_{i}$, a segment $s_{i}$ can be pruned if
$\theta^{c-j}\mathcal{D}_{s}(p_{o}^{j},\mathcal{M}(s_{i}))>\textit{dist}_{\textit{min}},$ (6)
@@
Given the current $k$-th smallest Object-Trajectory Real-time Distance $\textit{dist}_{k}$ between a moving object $o$ at timestamp $t$ with $o_{t}=(p_{o}^{c},T_{o})$ and the computed historical trajectories, a trajectory $T_{i}$ can be pruned if
$\mathcal{D}^{l}(p_{o}^{j},T^{\prime}_{i})\geq\frac{\textit{dist}_{k}-(1-\alpha)\mathcal{D}_{T}(p_{o}^{d},T_{i}^{L})}{\alpha\cdot\theta^{c-j}},$ (7)
@@
$\mathcal{D}^{l}(p_{o}^{j},T^{\prime}_{i})=\text{min}\{d(p_{o}^{j},p_{i})|p_{i}\in T^{\prime}_{i}\}.$ (8)
Lemma 2.
Given a moving object $o$ at timestamp $t$ with $o_{t}=(p_{o}^{c},T_{o})$ and a trajectory $T_{i}$, the Historical Trajectory Distance $\mathcal{D}_{H}(T_{o},T^{\prime}_{i})$ is no less than $d(p_{o}^{c},p_{i}^{x})$.
@@
According to Lemma 2, an early pruning strategy can be used before fully computing the distance between $o$ and $T_{i}$. $T_{i}$ can be pruned if $d(p_{o}^{c},p_{i}^{x})\geq(\textit{dist}_{k}-(1-\alpha)\cdot\mathcal{D}_{T}(p_{o}^{d},T_{i}^{L}))/\alpha$.
@@
Case 1: If the pivotal point ID in $T_{i}$ at $t$ is unchanged, then the TTD $\textit{dist}_{T}$ is unchanged and the HTD $\textit{dist}_{H}$ can be updated as follows:
$\textit{dist}_{H}=\text{max}\{\theta\cdot\textit{dist}^{\prime}_{H},\text{min}\{d(p_{o}^{c},p_{i})|p_{i}\in T^{\prime}_{i}\}\}.$ (9)
@@
Case 3: If the pivotal point ID in $T_{i}$ at $t$ is less than $x^{\prime}$, $\textit{dist}_{H}$, and $\textit{dist}_{T}$ can be computed according to Equations 2 and 3.
@@
However, this strategy may introduce errors to $\mathcal{D}_{H}(T_{o},T^{\prime}_{i})$ since some points in $T_{i}$ are ignored. We deduce the lower and upper error bounds as follows:
$\mathcal{E}^{l}=\mathcal{D}_{H}(T_{o},T^{\prime}_{i})-d(p_{o}^{c},p_{i}^{x}),$ (10)
$\mathcal{E}^{u}=\text{max}\{\theta^{c-j}\cdot d(p_{o},p_{i}^{j})|p_{o}\in T_{o},p_{i}^{j}\in T^{\prime}_{i}(g)\}-\mathcal{D}_{H}(T_{o},T^{\prime}_{i}).$ (11)
The lower error bound $\mathcal{E}^{l}$ can be easily proved according to Lemma 2. The upper error bound can be proved as follows.
@@
$\gamma=|x-x^{\prime}|\mbox{ mod }g.$ (13)
@@
IV-C OTRD Computation
The similarity computation between a moving object $o$ and a trajectory $T_{i}$ is performed in several key steps. First, the algorithm identifies the pivotal point $p_{i}^{x}$ of $T_{i}$ to the current position $p_{o}^{c}$ of $o$. Next, it computes the distance between $o$’ target destination $p_{o}^{d}$ and the line $L(p_{i}^{x},p_{i}^{n})$, where $p_{i}^{n}$ is the end point of $T_{i}$. Afterward, the distance between the $o$’s current trajectory $T_{o}$ and the sub-trajectory $T^{\prime}_{i}=\langle p_{i}^{1},...,p_{i}^{x}\rangle$ from $T_{i}$. To enhance the efficiency during this process, the speed-up strategies are employed to avoid unnecessary distance computations. Finally, the overall similarity is determined by summing these distances with a trade-off parameter $\alpha$, providing a comprehensive measure of how closely the moving object’s trajectory aligns with the historical one.
@@ IV-D CSTS Algorithm
IV-E Complexity Analysis
Complexity of OTRD Computation. Let $n_{o}$ and $n_{i}$ denote the numbers of points in the moving object trajectory $T_{o}$ and a historical trajectory $T_{i}$. Without optimization, the distance computation requires comparing all point pairs, giving a complexity of $\mathcal{O}(n_{o}\cdot n_{i})$. The OTRD_Computation function employs four optimizations: (1) Segment pruning reduces examined segments, effectively decreasing $n_{i}$. (2) $k$-bound pruning terminates early when the distance exceeds the $k$-th smallest value, reducing $n_{i}$ or $n_{o}$. (3) Buffer reuse avoids redundant computations for unchanged parts, further limiting $n_{i}$ or $n_{o}$. (4) Granularity adjustment samples points with step size $g$, reducing point-level comparisons to roughly $n_{i}/g$. With these, the average complexity becomes: $\mathcal{O}(n^{\prime}_{o}\cdot n^{\prime}_{i}/g)$, where $n^{\prime}_{o}$ and $n^{\prime}_{i}$ are the effective numbers of points after pruning and buffer reuse. In practice, $n^{\prime}_{o}\ll n_{o}$ and $n^{\prime}_{i}\ll n_{i}$, resulting in a substantial reduction in computation time.
Complexity of CSTS Algorithm. Let $|\mathcal{T}_{r}|$ be the number of candidate trajectories retrieved from the SVTI index at each timestamp. The range query cost is $\mathcal{O}(\log|\mathcal{T}|+|\mathcal{T}_{r}|)$, and each similarity computation costs $\mathcal{O}(n^{\prime}_{o}\cdot n^{\prime}_{i}/g)$. Maintaining the top-$k$ heap adds $\mathcal{O}(|\mathcal{T}_{r}|\log k)$. Thus, per-timestamp complexity is: $\mathcal{O}(\log|\mathcal{T}|+|\mathcal{T}_{r}|\cdot(n^{\prime}_{o}\cdot n^{\prime}_{i}/g+\log k))$. In practice, pruning and indexing greatly reduce $n^{\prime}_{i}$, $n^{\prime}_{o}$, $|\mathcal{T}_{r}|$, enabling efficient real-time search.
V Experiments
V-A Overall Experimental Settings
All experiments were conducted on a Linux server running Ubuntu with an Intel Xeon Icelake processor featuring 60 cores at 5.7 GHz and 240 GB of memory. The system is equipped with dual NVIDIA A10 GPUs, each providing 23 GB VRAM, for a total of 46 GB of GPU memory. All neural network models are implemented using PyTorch and trained on the GPU, and all index construction and search algorithms are implemented in Python (Code for CSTS [^1] https://github.com/hyLiu1994/ACTIVE (Last access: 2025.10)).
TABLE II: Dataset statistics.
@@ V-A Overall Experimental Settings
Datasets. We evaluate the proposed methods on the following datasets. Dataset statistics are reported in Table II.
  • RQ1. Evaluation on CSTS: How does the search algorithm CSTS perform in terms of efficiency and effectiveness?
  • RQ2. Ablation Study: What is the impact of the speed-up strategies on overall performance?
The evaluation on SVTI could be found in Appendix.
Metrics. The metrics for each task are listed in Table III. To evaluate the performance of CSTS, we select ten query instances, execute the selected queries ten times, and report the average. For effectiveness evaluation, we employ the hit rate ($\mathit{Hit}$):
$\mathit{Hit}=\frac{\lvert\mathcal{T}_{f}\cap\mathcal{T}\rvert}{k},$ (14)
TABLE III: Metrics and baselines.
@@
RQ1 Query Time (ms), Hit (%) O-PQT, O-PRT, H/F-TrajI, H/F-DFT, TrajCL-TI, CLEAR-TI
@@
Baselines. We compare the proposed methods with several baseline methods, as outlined in Table III. Specifically, CSTS is compared with six traditional methods and two learning-based methods to evaluate its efficiency and effectiveness. The baseline methods are listed as follows:
@@
  • H-TrajI/F-TrajI: Hausdorff/Fréchet + Traj_Index [8]. An R-Tree-based similar trajectory method using Hausdorff or Fréchet distance, where the index is built according to the centroid of each data trajectory’s MBR.
  • H-DFT/F-DFT: Hausdorff/Fréchet + DFT [8]. A Distributed Framework for Trajectory similarity search (DFT) using Hausdorff or Fréchet distance.
  • TrajCL-TI: TrajCL [28] + SVTI. A trajectory similarity search method based on SVTI, employing a graph-based long-term dependency model for similarity computation.
  • CLEAR-TI: CLEAR [29] + SVTI. A trajectory similarity search method using SVTI with a similarity computation technique based on contrastive representation learning.
@@
TABLE IV: Parameter settings.
  Parameters Questions Settings
$\Lambda$ RQ1 D-1, D-2, D-3, D-4, D-5
$l_{q}$ RQ1 10, 20, 30, 40, 50
$r$ RQ1 5, 10, 15, 20, 25
$k$ RQ1 30, 40, 50, 60, 70
$\theta$ RQ1 0.45, 0.5, 0.55, 0.6, 0.65
$\alpha$ RQ1 0.45, 0.5, 0.55, 0.6, 0.65
$g$ RQ2 1, 5, 10, 15, 20
Parameters. The parameter settings are summarized in Table IV, with default values in bold. A description of each parameter is provided below:
@@
V-B Evaluation on CSTS